How much money should Sheila invest in an account that earns 8% annual interest compounded monthly in order to have $4,000 in 5 years?
Answer by rfer(16322)

P=4000/(1+.08/12)^12*5
P=4000/(1.006666)^60
P=4000/1.4897865
P=$2685